OpenStudy (anonymous):

How do you factor this quadratic equation? 6(x^2-4x+4)^2 + (x^2-4x+4) -1

OpenStudy (anonymous):

\[6\left( x ^{2}-4x+4 \right)^{2}+ (x ^{2}-4x+4)-1\]

OpenStudy (anonymous):

Is that the final answer?

OpenStudy (anonymous):

\[6(x^{2}-4x+4)^{2}+(x^{2}4x+4)-1\\ = 2(x^{2}-4x+4)+1)(3)(x^{2}-4x+4)-1\\ (2x^{2}-8x+9)(3x^{2}-12x+12)-1\\ (2x^{2}-8x+9)(3x^{2}-12x+11) \]